OpenStudy (anonymous):

What is the measure of the complement of an angle that measures 21°?

OpenStudy (anonymous):

@iambatman

jimthompson5910 (jim_thompson5910):

x = unknown angle x+21 = 90 solve for x

jimthompson5910 (jim_thompson5910):

the angles x and 21 are complementary, so they form a right angle (ie they add to 90 degrees)

OpenStudy (anonymous):

69

jimthompson5910 (jim_thompson5910):

correct, check: x+21 = 90 69+21 = 90 90 = 90 ... works out

OpenStudy (anonymous):

Thanks